![]() |
![]() |
![]() |
![]() |
(Week 11, 2003, 10 Mar. - 16 Mar.)
In the display above (click to enlarge), the solar X-ray fluence seen by RESIK (black points) is plotted atop GOES red line. Only periods with good spectral measurements are indicated. The total number of good spectral measurements is given in the title line.
Operation mode: |
2
s DGI has been used.
|
|
Calibrations made: |
No calibrations made
| |
Switch to backup power supply unit: |
As explained below, on 14 March, we sent command to switch RESIK power supply to redundant unit. | |
ADS Settings: |
||
|
(ORD1) HV: Det. A - 1450 V, Det. B - 1419 V. | (ORD3) HV: Det. A - 1389 V, Det. B - 1328 V. | ||||||
Channel |
l Band |
ADS |
Channel |
l Band |
ADS | ||
#1 | #2 | 3.37 - 3.88 Å | 55 - 110 | #1 | #2 | 1.16 - 1.29 Å | 110 - 230 |
#2 | #0 | 3.82 - 4.33 Å | 80 - 165 | #2 | #0 | 1.27 - 1.44 Å | 110 - 230 |
#3 | #3 | 4.31 - 4.89 Å | 90 - 160 | #3 | #3 | 1.44 - 1.63 Å | 110 - 230 |
#4 | #1 | 4.96 - 6.09 Å | 135 - 200 | #4 | #1 | 1.65 - 2.03 Å | 140 - 230 |
|
![]() |
Data gaps due to missing telemetry: ~ 24.3 h.
Note, that the link to the catalogue page is now: http://www.cbk.pan.wroc.pl/resik_catalogue.htm.
RESIK Switched to Backup Power Unit
As announced recently, (weekly 10, 2003), we noted fast changes of the fluorescence peak position on all PHA channels. This finding finally assured us that the reason of recent disturbances in RESIK operations (frequent single event upsets, switches to backup processors) was related to ailing power supply unit. On March 14 at 08:27 UT, the command has been sent to CORONAS-F S/C, switching the power supply to backup unit. After switch-up, RESIK resumed operations with somewhat changed characteristics. In the Figure, we show the plot of fluorescence peak position during the transition to backup.
Figure 1. Plot showing recent changes of the fluorescence peak position as observed for detectors A and B.
These plots show that the HV became stable again, however,
the positions of fluorescence peak moved to lower PHA bins in all channels. We interpret this as indication of the
return of HV levels to the initial settings, these not altered by ailing power supply. The observed changes are by
0.45, 0.60, 0.15 and 0.25 PHA bins for channels (software) #0, #1, #2 and #3 respectively. This change, although
relatively small, might imply small change of the ADS (amplitude discriminator settings). In this
circumstances, new in-flight ADS calibration will be commanded as soon as
possible.
Prepared by: Janusz Sylwester js@cbk.pan.wroc.pl and the RESIK engineering team: Mirek Kowalinski mk@cbk.pan.wroc.pl and Witold Trzebinski wt@cbk.pan.wroc.pl.
The report presented has been done in "real time" and so it may contain jargon, blunders, or trivialities. We do not have also an English native speaker in our Wroclaw group! We would be happy to discuss problems mentioned above in more details if necessary.
RESIK data are in the open public domain and
can be requested from: http://surfwww.mssl.ucl.ac.uk/surf/data_request.html.
Previous RESIK_weekly notes are in the archive: http://www.cbk.pan.wroc.pl/resik_archive.htm
Page made on 21 March 2003 by: Anna Kepa ak@cbk.pan.wroc.pl and Jarek Bakala jb@cbk.pan.wroc.pl